OpenStudy (anonymous):

Please help. ln |cos 0| - ln |sin 0|

OpenStudy (anonymous):

\[\sin(0)=0\] so \[\ln(\sin(0))\] is undefined

OpenStudy (anonymous):

thats none of the answers

OpenStudy (unklerhaukus):

what are your options/

OpenStudy (anonymous):

it might be none of the answers, but you can take it to the bank that there is no log of zero.

OpenStudy (anonymous):

ln |cot0| ln |tan0| ln| sec0| ln|cos0| ln|sin0| ln|csc0|

OpenStudy (anonymous):

oh perhaps that is a \[\theta\]??

OpenStudy (anonymous):

yeah its theta

OpenStudy (anonymous):

in which case the answer is \[\ln(\frac{\cos(\theta)}{\sin(\theta)})\] \[\ln(\cot(\theta)\]

OpenStudy (anonymous):

in general \[\theta\neq 0\]

OpenStudy (unklerhaukus):

\[ln |cos \theta | - ln |sin \theta|\]\[=ln \big|{cos \theta \over sin\theta }\big |\]\[=ln \big|{cot \theta }\big |\]

OpenStudy (anonymous):

that was perfect